OpenStudy (anonymous):

given y varies directly as x, and y=9 when x=3. Find y when x=11

OpenStudy (anonymous):

it would be 17

OpenStudy (anonymous):

y = ax 9 = a3 so a = 3 y = 3x y = (3)(11)

OpenStudy (anonymous):

So, y would equal 33.

OpenStudy (anonymous):

no.... it wouldn't

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @lvasquev462

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

i have one more of these questions

OpenStudy (anonymous):

sure :P

OpenStudy (anonymous):

go right ahead, np.

OpenStudy (anonymous):

given y varies directly as x and y=4 when x=20. Find x when y=6

OpenStudy (anonymous):

y = ax 4 = a(20) a = (1/5) y = (1/5)x x = 5y x = (5)(6) = 30

OpenStudy (anonymous):

y=ax y=(20)(6)

OpenStudy (anonymous):

it'S 20 AND 6! URG

OpenStudy (anonymous):

So, all I did was basically the same thing, but the only twist here is that I flipped the "x" and "y". See that?

OpenStudy (anonymous):

yeah i seen tht

OpenStudy (anonymous):

So, with these "varying directly" problems, you can start out the same way in both cases. Nice working with you!